caira12:

Ava, Grace, Jayden, and Zoe together stacked a total of 50 brownies on a tray. Ava stacked 0.14 of the brownies, Grace stacked 4 of the 50 brownies, Jayden stacked 24% of the brownies, and Zoe stacked the remaining brownies. Who stacked the most brownies? (5 points) Select one: a. Ava b. Grace c. Jayden d. Zoe

Mercury:

well first you need to figure out: grace: 0.14 of 50 grace: 4/50 jayden: 24% of 50 zoe: 50 - every body else's # so you have to calculate each one and see which one's the largest let me know if you're getting stuck or want to check your sol'n

Mercury:

well, how about we break it down step by step "grace: 0.14 of 50" "of" usually means multiply in math, so simply multiply 0.14 * 50 to see how many brownies grace stacked

caira12:

7

Mercury:

good! now let's do Jayden's what is 24% of 50?

NonportritV2:

Here's the answer: 50 - (50 x 0.14) - 4 - (50 x 24%) = 50 - 7 - 4 - 12 = 50 - 23 = 27 Therefore Zoe stacked the most brownies. She stacked 27 brownies out of the 50 brownies. The answer is option D "Zoe." Good luck.

caira12:

12

Mercury:

good! so now you simply subtract 50 - 4 - 7 - 12 = ? to get zoe's count

caira12:

thank you